OpenStudy (anonymous):

Write x^2 - 8x + 13 = 0 in the form (x - a)2 = b, where a and b are integers.

OpenStudy (anonymous):

(x - 4)^2 = 3 (x - 3)^2 = 2 (x - 2)^2 = 1 (x - 1)^2 = 4

OpenStudy (anonymous):

@Ashleyisakitty @SolomonZelman